OpenStudy (stavii):

What is the value of a? 9a-3(a-6) =-6 a=________?

OpenStudy (king.void.):

Welcome to openstudy!! Wrong section ;) please close this and post it in the correct section. Thanks! Void!~

OpenStudy (iloveopenstudy2):

9a-3(a-6) =-6 First let's do this part: 3(a-6) Use the distributive property a(b-c) = ab - ac So what is 3(a-6) = ??

OpenStudy (anonymous):

9a-3(a-6)=-6 9a-3a-18=-6 9a-3a=12 6a=12 a=2 Hope this helps! Enjoy your day!

OpenStudy (confluxepic):

Welcome to OpenStudy! @STAVII

OpenStudy (stavii):

Thank-you guys that was very helpful!